OpenStudy (anonymous):

Simplify the expression. (6 - )(6 +

OpenStudy (anonymous):

\[(6-\sqrt{11)}(6+\sqrt{11})\] how od you solve this

OpenStudy (anonymous):

FOIL method

OpenStudy (anonymous):

n whats and how do you do that

OpenStudy (anonymous):

First Outter Inner Last

OpenStudy (anonymous):

huh i suck at math

OpenStudy (anonymous):

multiple the first of each parentheses then the inners then outer then lasts

OpenStudy (anonymous):

36 and 121

OpenStudy (anonymous):

can you show me the steps plz

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so after you FOIL it out.... you get 36+6√11-6√11-√121

OpenStudy (anonymous):

got that?

OpenStudy (anonymous):

√121 turns into 11 and the 6√11 they both cancel out because one is positive and ones negative. then you end up with 36-11 which = 25

OpenStudy (anonymous):

thank you